OpenStudy (anonymous):

the lengths of the sides of a triangle are consecutive odd integers. if the perimeter is 1 less than four times the shortest side, find the lenght of each side

jimthompson5910 (jim_thompson5910):

(2x+1)+(2x+3)+(2x+5) = 4(2x+1)-1 2x+1+2x+3+2x+5 = 8x+4-1 6x+9=8x+4-1 6x+9=8x+3 6x=8x+3-9 6x-8x=3-9 -2x=3-9 -2x=-6 x=-6/(-2) x=3 So the three side lengths are 2(3)+1 = 6+1 = 7 2(3)+3 = 6+3 = 9 2(3)+5 = 6+5 = 11

OpenStudy (anonymous):

thanks but how did you get 2x

OpenStudy (anonymous):

nevermind i got it